2 NH3  +  3 CuO → 3 Cu  +  N2  +   3 H2O

In the above equation, how many grams of N2 can be made when 18.4 moles of CuO are consumed?

Round your answer to the nearest tenth. If you answer is a whole number like 4, report the answer as 4.0

Use the following molar masses. If you do not use these masses, the computer will mark your answer incorrect.:


1
Expert's answer
2022-02-06T11:19:53-0500

The molar mass of CuO is 63.5 + 16 = 79.5

The molar mass of N2 is 2×14 = 28


According to the equation, for every 3 molecules of CuO, one molecule of N2 is made.

Therefore 3×79.5 g of CuO make 28 grams of N2.

Therefore 195 g of CuO make 28×(195/3×79.5) = 22.9 g of N2.